(3) Use both methods to solve the equation. 2 – 4 – 2g = tetparameters and undetermimed coefficiemts
Answer #1

y - y - 2y = id edit Auxiliary equation. me-m-2=0 = (m-2) (M +1) = 0 > m=-1,2 C.F. - Ciert + C2 det Salve with undetermined